# Inverse Virtual Try-On: Generating Multi-Category Product-Style Images from Clothed Individuals

### v1

@@ 3.2 Text-Enhanced Garment Try-Off
Multimodal Hybrid Attention. For the first problem, we assume that it can be addressed by captions focused on structural details and hidden parts. To validate our claim, we propose a new MHA module to seamlessly mix text information, intermediate features from $F_{E}$ and latent features of the denoising DiT. Inspired by the key findings in SD3 esser2024scaling , we concatenate the text features with the visual inputs along the sequence length dimension, thus obtaining:
@@
Text embeddings are constructed via the concatenation of CLIP radford2021learning ***As previously mentioned, we consider the combined embedding from CLIP ViT-L and Open-CLIP bigG/14. and T5 raffel2020exploring encoders applied to the input caption $c$ as follows:
$\bm{e}_{\text{text}}=[\operatorname{CLIP}(c),\operatorname{T5}(c)],\quad\text{with }\bm{e}_{\text{text}}\in\mathbb{R}^{77\times 4096}.$ (6)
Visual Category Conditioning with Text Modulation. While some datasets are restricted to upper-body garments choi2021vitonhd , others, such as Dress Code morelli2022dresscode , encompass a broader range, including lower-body items and full-body dresses. This variability introduces ambiguity in garment structure and scale, motivating the need for an explicit encoding of category-level priors.
To address this, we employ high-level conditioning via AdaLN modulation huang2017arbitrary . As shown in previous works garibi2025tokenverse , these layers can be successfully leveraged to adapt “appearance” or “style” information into existing Transformer-based architectures. For this reason, we extract a pooled textual representation $\bm{e}_{\text{pooled}}\in\mathbb{R}^{2048}$ of CLIP textual features of the caption $c$ and inject them into the model through the modulation layers, following Eq. 4. The pooled vector $\bm{e}_{\text{pooled}}\in\mathbb{R}^{2048}$ encapsulates a coarser representation than the full textual embeddings $\bm{e}_{\text{text}}\in\mathbb{R}^{77\times 4096}$, thus being suitable for high-level information conditioning.
We generate textual descriptions using Qwen2.5-VL bai2025qwen2 in a zero-shot setting. We select this model due to its state-of-the-art performance among open-source multimodal large language models. To align with our objective of capturing fine-grained structural attributes, we steer the captioning process to emphasize garment-level semantics, such as garment type, cut, or sleeve length, while deliberately omitting low-level visual features (e.g., color or texture). Empirically, we find that our proposed MHA module is sufficient for accurate color transfer without explicit textual supervision, allowing us to focus the text modality on structural conditioning.
We train the main DiT module (i.e., $F_{D}$) following a diffusion loss with multiple conditioning signals:
$\mathcal{L}_{\text{DiT}}=\mathbb{E}_{\bm{z}_{g},\epsilon_{t},t}\left[\left\|\bm{z}_{g}-F_{D}(\bm{z}_{t},\bm{e}_{\text{pooled}},F_{E}(\bm{z}_{0}^{\prime},\bm{x}_{\text{model}},0),t)\right\|^{2}\right],$ (7)
with $F_{E}(\bm{z}_{0}^{\prime},\bm{x}_{\text{model}},0)$ being the list of keys and values extracted from $F_{E}$ at timestep $t=0$. We extract this list from $F_{E}$ at $t=0$ and re-use them in $F_{D}$ for all subsequent timesteps, as we want to use key/values from clean data.
@@ 3.3 Garment Aligner
While our model is effective at generating realistic and structurally coherent garments, we observe occasional failures in preserving high-frequency details such as fine-grained textures and logos. We hypothesize two primary contributing factors: (i) the diffusion loss $\mathcal{L}_{\text{diff}}$, defined in the noise space, optimizes over perturbed latents rather than directly over image-space reconstructions, limiting its sensitivity to fine-grained patterns; and (ii) the inherent generation dynamics of diffusion models, where errors introduced in early timesteps – typically encoding low-frequency content – can accumulate and degrade the fidelity of high-frequency details in later stages. To mitigate this, we draw inspiration from REPA yu2025representationalignmentgenerationtraining , and propose to explicitly align the internal feature representation of our DiT with that of a pre-trained vision encoder. Specifically, we encourage patch-wise consistency between the eighth Transformer block of our main DiT model $F_{D}$ and the corresponding features extracted from DINOv2 oquab2024dinov2learningrobustvisual .
Let $\bm{h}_{\text{DiT}}\in\mathbb{R}^{3072\times d}$ denote the token sequence obtained from the eighth Transformer block of the DiT decoder $F_{D}$, corresponding to a $64\times 48$ patch grid with embedding dimension $d$. Separately, let $\bm{h}_{\text{enc}}\in\mathbb{R}^{1024\times d^{\prime}}$ be the $32\times 32$ token grid extracted from a frozen DINOv2 encoder, with embedding dimension $d^{\prime}$ (where $d^{\prime}\neq d$). To bridge this mismatch, we introduce a lightweight garment aligner module composed of a convolutional neural network $\phi_{\text{CNN}}:\mathbb{R}^{64\times 48\times d}\rightarrow\mathbb{R}^{32\times 32\times d^{\prime}}$ which is used to downsample the spatial token grid while preserving local structure and to project the token embeddings into the DINOv2 feature space. The aligned tokens are defined as $\tilde{\bm{h}}_{\text{DiT}}=\phi_{\text{CNN}}(\bm{h}_{\text{DiT}})\in\mathbb{R}^{1024\times d^{\prime}}$.
@@
where $\tilde{\bm{h}}_{i}^{\text{DiT}}$ and $\bm{h}_{i}^{\text{enc}}$ are the $i$-th aligned and reference tokens, respectively, $i$ is the patch index, $N$ is the total number of tokens, and $\cos$ is the cosine similarity.
Overall Loss Function. Our final training objective combines the standard diffusion loss $\mathcal{L}_{\text{diff}}$ with the garment alignment loss $\mathcal{L}_{\text{align}}$ previously introduced. The overall objective is thus defined as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{diff}}+\lambda\cdot\mathcal{L}_{\text{align}},$ (9)
@@ 4 Experiments
4.1 Experimental Settings and Datasets
We conduct our experiments using two publicly available fashion datasets: VITON-HD choi2021vitonhd and Dress Code morelli2022dresscode . Both datasets provide images at a resolution of $1024\times 768$. VITON-HD contains only upper-body garments and represents a single-category setting, while Dress Code includes multiple categories (i.e., dresses, upper-body, and lower-body garments) enabling evaluation of the generalization capabilities of our methods across diverse garment types. For both the feature extractor and the diffusion backbone, we adopt Stable Diffusion 3 medium esser2024scaling . All models are trained on a single node equipped with 4 NVIDIA A100 GPUs (64GB each), using DeepSpeed ZeRO-2 rajbhandari2020zeromemoryoptimizationstraining for efficient distributed training. We use a total batch size of 32 and train each model for 30k steps, corresponding to approximately 960k images. Optimization is performed with AdamW loshchilov2019decoupledweightdecayregularization , using a learning rate of $1\times 10^{-4}$, a warmup phase of 3k steps, and a cosine annealing schedule. We train separate models per dataset to account for differences in distribution and garment structure. In all experiments, we set the alignment loss weight $\lambda$ equal to $0.5$.
@@ 4.1 Experimental Settings and Datasets
Method SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$ SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$
TryOffDiff velioglu2024tryoffdiff 75.79 39.70 29.88 70.02 42.80 76.59 40.62 29.04 37.97 17.30
Any2AnyTryon guo2025any2anytryonleveragingadaptiveposition 77.56 35.17 25.17 12.32 3.65 76.61 38.99 25.78 15.77 3.22
TEMU-VTOFF (Ours) 75.95 31.46 18.66 5.74 0.65 74.54 35.48 19.75 10.94 0.76
@@
Method SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$ SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$
TryOffDiff velioglu2024tryoffdiff 74.10 43.35 32.54 162.02 137.10 76.79 35.29 27.88 72.19 47.58
Any2AnyTryon guo2025any2anytryonleveragingadaptiveposition 78.15 34.72 25.87 30.06 12.01 77.93 31.80 23.86 19.20 6.27
TEMU-VTOFF (Ours) 73.94 34.60 19.57 13.83 2.04 79.39 24.32 16.67 11.29 0.59
@@ 4.2 Comparison with the State of the Art
To evaluate the proposed TEMU-VTOFF architecture, we use a combination of perceptual, structural, and distributional similarity metrics. Specifically, we report LPIPS zhang2018perceptual , SSIM wang2004ssim , DISTS Ding_2020 , FID parmar2021cleanfid , and KID bińkowski2021demystifyingmmdgans . We compare our approach against recent VTOFF methods, including TryOffDiff velioglu2024tryoffdiff , TryOffAnyone xarchakos2024tryoffanyone , and Any2AnyTryon guo2025any2anytryonleveragingadaptiveposition , the latter being a more general framework designed for both VTOFF and VTON tasks. Notably, TryOffDiff and TryOffAnyone are trained exclusively on VITON-HD, which contains only upper-body garments. In contrast, Any2AnyTryon is trained on a mixture of datasets including Dress Code, VITON-HD, and DeepFashion2 ge2019deepfashion2versatilebenchmarkdetection , enabling broader category coverage.
For fair comparison on the multi-category Dress Code dataset, we retrain TryOffDiff from scratch using its publicly available code and default settings. Due to the lack of released training code, we are unable to retrain TryOffAnyone. This setup allows us to properly evaluate how well existing methods generalize across multiple categories and to highlight the robustness and flexibility of our approach in handling different garment types.
@@
Results on the Dress Code Dataset. Table 1 presents the image quality metrics on the Dress Code dataset. Our method significantly outperforms existing state-of-the-art approaches across all evaluation metrics. The results are especially pronounced in this dataset, with respect to prior works, because of its multi-category setting. This shows that our method is category-agnostic and benefits from the joint use of textual garment descriptions and fine-grained masks. As a result, our model achieves a better perceptual quality and better alignment with the ground-truth distribution.
@@
In Fig. 3, we provide qualitative results comparing TEMU-VTOFF with competitors. This shows the challenges introduced by the Dress Code different set of categories. TryOffDiff often fails to preserve key visual characteristics such as color, texture and shape. Furthermore, it struggles with the reconstruction of lower-body items. Instead, Any2AnyTryon performs better, but often fails with the generation of the target garment. In contrast, our method is able to closely match the target garment across all categories.
Results on the VITON-HD Dataset. In Table 2, we report the quantitative results on the VITON-HD dataset. Also in this setting, TEMU-VTOFF achieves consistent improvements over existing approaches across all metrics. However, the improvements are less pronounced compared to those observed on the Dress Code dataset. This discrepancy is primarily due to the more constrained nature of VITON-HD, which focuses exclusively on upper-body garments. By contrast, Dress Code includes different categories, including full-body and lower-body items like dresses, skirts, and pants, which are more prone to occlusions and complex geometries. In these cases, the use of textual descriptions and fine-grained masks becomes crucial, offering strong guidance that significantly enhances garment reconstruction quality. Thus, the improvements of our approach are especially evident in more challenging multi-category scenarios.
A visual comparison on sample VITON-HD images is shown in Fig. 4, which further demonstrates the improved garment reconstruction quality of our proposed method.
@@
Table 2: Quantitative results on the VITON-HD dataset. $\uparrow$ indicates higher is better, $\downarrow$ lower is better.
Method SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$
TryOffDiff velioglu2024tryoffdiff 75.53 39.56 25.53 17.49 5.30
TryOffAnyone xarchakos2024tryoffanyone 75.90 35.26 23.47 12.74 2.85
Any2AnyTryon guo2025any2anytryonleveragingadaptiveposition 75.72 37.95 24.32 12.88 3.01
TEMU-VTOFF (Ours) 77.21 28.44 18.04 8.71 1.11
@@ 4.3 Ablation Studies
To assess the contribution of each component in our pipeline, we conduct a detailed ablation study on the Dress Code dataset, as shown in Table 3. Specifically, we evaluate the impact of removing the textual description of the garment, the fine-grained mask, or both. Further, we investigate the impact of our garment aligner. The results show that all the components play a crucial role, demonstrating their contribution to improving the final image quality.
@@
SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ DISTS $\downarrow$ FID $\downarrow$
@@
w/o text and masks 71.04 39.68 25.20 9.63 3.17 23.71 19.75 65.85 49.19 20.12 15.47
w/o text modulation 73.88 34.63 22.54 7.75 1.52 24.02 13.48 24.33 18.13 19.27 13.30
w/o fine-grained masks 74.65 32.33 20.87 6.58 1.03 20.85 11.31 22.34 15.74 19.42 13.62
TEMU-VTOFF (Ours) 75.95 31.46 18.66 5.74 0.65 19.75 10.94 19.57 13.83 16.67 11.29
Effect of Garment Aligner Component
w/o garment aligner 76.01 30.84 20.63 5.91 0.78 21.77 11.26 22.26 14.22 17.86 11.86
TEMU-VTOFF (Ours) 75.95 31.46 18.66 5.74 0.65 19.75 10.94 19.57 13.83 16.67 11.29
@@
Figure 5: Qualitative comparisons validating the effectiveness of the proposed components on the Dress Code dataset.
To better understand the strength of each component proposed with our approach, in Fig. 5 we provide a visual comparison on the Dress Code dataset. Our method without textual conditioning relies exclusively on visual features extracted from the person wearing the garment, without any textual guidance. Notably, this often results in failures in challenging cases, especially for garments with strong occlusions such as bodysuits or garments partially hidden by arms or hair. This limitation has been addressed by conditioning the generation to also use a textual structural description of the target garment. This structural guidance enables the model to better capture the structure of the garment, leading to significantly more accurate outputs in occluded scenarios. In addition, using fine-grained masks and the garment aligner further improves the quality of the generated garments.
Limitations. While our method shows strong performance and generalization, some limitations remain. First, it struggles to reconstruct fine-grained details such as logos or printed text – partly due to the SD3 medium backbone, which is known to handle such elements inconsistently. Second, performance on lower-body garments is less reliable than for upper-body garments and dresses, likely due to class imbalance in the Dress Code dataset. Additional discussion and failure cases are included in the supplementary material.

### v2

@@ 3.2 Dual-DiT Text-Enhanced Garment Try-Off
Multimodal Hybrid Attention. Our new module seamlessly mix text information, latent features of the denoising DiT, and intermediate features from $F_{E}$. Inspired by the key findings in SD3 (Esser et al., 2024), we concatenate the text features with the visual inputs along the sequence length dimension, thus obtaining:
@@
Text embeddings are constructed via the concatenation of CLIP (Radford et al., 2021) [^1] Following SD3, we consider the combined embedding from CLIP ViT-L and Open-CLIP bigG/14. and T5 (Raffel et al., 2020) encoders applied to the input caption $c$ as follows:
$\bm{e}_{\text{text}}=[\text{CLIP}(c),\text{T5}(c)],\quad\text{with }\bm{e}_{\text{text}}\in\mathbb{R}^{77\times 4096}.$ (5)
Now we pose a relevant question: is it possible to disambiguate the garment category from the mask alone? A mask input can improve multi-category handling by acting as a hard discriminator between two garments, in contrast to text, which acts as a soft discriminator since it does not directly indicate the pixels occupied by the target garment. Therefore, the mask can help to visually force the model to retain only upper- or lower-body information but it can not tell much about the appearance of a garment, because it is highly warped together with the person, resulting in visual artifacts. Textual information is critical, together with mask information, to extract the category information of the garment. To address this, we decide to use also the global conditioning scheme provided by AdaLN (Huang and Belongie, 2017) in SD3. As shown in previous works (Garibi et al., 2025), these layers can be successfully leveraged to adapt “appearance” or “style” information into existing Transformer-based architectures. For this reason, we extract a pooled textual representation $\bm{e}_{\text{pool}}\in\mathbb{R}^{2048}$ of CLIP textual features of the caption $c$ and inject them into the model through the modulation layers, following Eq. 3. The pooled vector $\bm{e}_{\text{pool}}\in\mathbb{R}^{2048}$ encapsulates a coarser representation than the full textual embeddings $\bm{e}_{\text{text}}\in\mathbb{R}^{77\times 4096}$, thus being suitable for high-level information conditioning.
Training. We employ a two-stage training procedure: we train the module $F_{E}$ alone, detached from the dual DiT $F_{D}$, according to the diffusion loss $L_{\text{diff}}$ defined as follows:
Then, we train the dual DiT $F_{D}$ following a diffusion loss with multiple conditioning signals:
$\mathcal{L}_{\text{DiT}}=\mathbb{E}_{\bm{z}_{g},\epsilon_{t},t}\left[\left\|\bm{z}_{g}-F_{D}(\bm{z}_{t},\bm{e}_{\text{pool}},F_{E}(\bm{z}_{0}^{\prime},\bm{x}_{\text{model}},0),t)\right\|^{2}\right],$ (7)
where $z_{g}=\mathcal{E}(x_{g})$ is the latent representation of the target garment encoded by the VAE, and with $F_{E}(\bm{z}_{0}^{\prime},\bm{x}_{\text{model}},0)$ being the list of keys and values extracted from $F_{E}$ at timestep $t=0$. We extract this list from $F_{E}$ at $t=0$ and re-use them in $F_{D}$ for all subsequent timesteps, as we want to use key/values from clean data.
@@ 3.3 Garment Aligner
While our model is effective at generating realistic and structurally coherent garments, we observe occasional failures in preserving high-frequency details such as fine-grained textures and logos. We hypothesize two primary contributing factors: (i) the diffusion loss $\mathcal{L}_{\text{diff}}$, defined in the noise space, optimizes over perturbed latents rather than directly over image-space reconstructions, limiting its sensitivity to fine-grained patterns; and (ii) the inherent generation dynamics of diffusion models, where errors introduced in early timesteps – typically encoding low-frequency content – can accumulate and degrade the fidelity of high-frequency details in later stages. To mitigate this, we draw inspiration from REPA (Yu et al., 2025), and propose to explicitly align the internal feature representation of our DiT with that of a pre-trained vision encoder. Specifically, we encourage patch-wise consistency between the eighth Transformer block of our main DiT model $F_{D}$ and the corresponding features extracted from DINOv2 (Oquab et al., 2023).
Formally, let $\bm{h}_{\text{DiT}}\in\mathbb{R}^{3072\times d}$ denote the token sequence obtained from the eighth Transformer block of the DiT decoder $F_{D}$, corresponding to a $64\times 48$ patch grid with embedding dimension $d$. Separately, let $\bm{h}_{\text{enc}}\in\mathbb{R}^{1024\times d^{\prime}}$ be the $32\times 32$ token grid extracted from a frozen DINOv2 encoder, with embedding dimension $d^{\prime}$ (where $d^{\prime}\neq d$). To bridge this mismatch, we introduce a lightweight garment aligner module composed of a convolutional neural network $\phi_{\text{CNN}}:\mathbb{R}^{64\times 48\times d}\rightarrow\mathbb{R}^{32\times 32\times d^{\prime}}$ which is used to downsample the spatial token grid while preserving local structure and to project the token embeddings into the DINOv2 feature space. The aligned tokens are defined as $\tilde{\bm{h}}_{\text{DiT}}=\phi_{\text{CNN}}(\bm{h}_{\text{DiT}})\in\mathbb{R}^{1024\times d^{\prime}}$.
@@
where $\tilde{\bm{h}}_{i}^{\text{DiT}}$ and $\bm{h}_{i}^{\text{enc}}$ are the $i$-th aligned and reference tokens, respectively, $i$ is the patch index, $N$ is the total number of tokens, and $\cos$ is the cosine similarity. It is important to note that the garment aligner is strictly a training-time component used to compute $\mathcal{L}_{align}$. It is discarded during inference, adding no computational overhead to the generation process.
Overall Loss Function. The garment aligner is applied in the second stage of our training. Our final training objective combines the standard diffusion loss $\mathcal{L}_{\text{DiT}}$ with the garment alignment loss $\mathcal{L}_{\text{align}}$ previously introduced. The overall objective is thus defined as:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{DiT}}+\lambda\cdot\mathcal{L}_{\text{align}},$ (9)
@@ 4 Experiments
Method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$
TryOffDiff - - - - - - 76.59 11.54 40.62 29.04 37.97 17.30
Any2AnyTryon 77.56 12.67 35.17 25.17 12.32 3.65 76.61 12.27 38.99 25.78 15.77 3.22
MGT 77.77 11.99 35.37 27.28 13.47 5.28 76.77 11.44 39.70 28.13 19.49 6.87
TEMU-VTOFF 75.95 12.90 31.46 18.66 5.74 0.65 74.54 12.51 35.48 19.75 10.94 0.76
@@
Method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$
Any2AnyTryon 78.15 12.42 34.72 25.87 30.06 12.01 77.93 13.32 31.80 23.86 19.20 6.27
MGT 77.29 11.64 36.31 28.00 25.98 9.64 79.26 13.09 30.11 25.70 19.09 5.74
TEMU-VTOFF 73.94 12.14 34.60 19.57 13.83 2.04 79.39 14.36 24.32 16.67 11.29 0.59
@@ 4.1 Comparison with the State of the Art
We conduct our experiments using two publicly available fashion datasets: VITON-HD (Choi et al., 2021) and Dress Code (Morelli et al., 2022). VITON-HD contains only upper-body garments and represents a single-category setting, while Dress Code includes multiple categories (i.e., dresses, upper-body, and lower-body garments) enabling evaluation of the generalization capabilities of our methods across diverse garment types. To evaluate the proposed TEMU-VTOFF architecture, we use a combination of perceptual, structural, and distributional similarity metrics. Specifically, as VTOFF is a paired generation setting, we mainly rely on reference-based like SSIM (Wang et al., 2004), PSNR (Wang et al., 2004), LPIPS (Zhang et al., 2018), and DISTS (Ding et al., 2020), alongside FID (Parmar et al., 2022) and KID (bińkowski2021demystifyingmmdgans). We compare our approach against recent VTOFF methods, including TryOffDiff (Velioglu et al., 2024), TryOffAnyone (Xarchakos and Koukopoulos, 2024), MGT (Velioglu et al., 2025), Voost (Lee and Kwak, 2025), One Model For All (Liu et al., 2025), and Any2AnyTryon (Guo et al., 2025). For TryOffAnyone, Voost, and One Model For All, we report the results only on the VITON-HD dataset because they have not been trained on the Dress Code dataset. Additionally, we retrain TryOffDiff on Dress Code using the official code and hyperparameters provided by the authors. Since TryOffDiff is not designed to handle multi-category garments, we report results only for the upper-body category.
@@
Results on the Dress Code Dataset. Table 1 reports the experimental results on the Dress Code dataset. As observed, our method outperforms existing state-of-the-art approaches across most evaluation metrics and garment categories. These results indicate that our approach is category-agnostic and benefits from the joint use of textual garment descriptions and fine-grained masks. Consequently, our model achieves a better perceptual quality and closer alignment with the ground-truth distribution compared to competing methods. Performance on lower-body garments is slightly lower due to the class imbalance in the Dress Code dataset, which contains significantly fewer lower-body samples ($\sim$9k) compared to upper-body ($\sim$15k) and full-body dresses ($\sim$29k).
@@
In Fig. 3, we provide qualitative results comparing TEMU-VTOFF with competitors. These examples highlight the challenges posed by the diverse set of categories in Dress Code. As shown, MGT and Any2AnyTryon frequently struggle to preserve key visual attributes such as color, texture, and shape. In contrast, our method is able to closely match the target garment across all categories, demonstrating a clear improvement in generation quality.
@@
Table 2: Quantitative results on the VITON-HD dataset. $\uparrow$ indicates higher is better, $\downarrow$ lower is better. $\dagger$ denotes results taken directly from the original papers.
Method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$
TryOffDiff 75.53 11.65 39.56 25.53 17.49 5.30
TryOffAnyone 75.90 12.00 35.26 23.47 12.74 2.85
Any2AnyTryon 75.72 12.00 37.95 24.32 12.88 3.01
MGT 78.10 - 36.30 24.70 21.90 8.90
Voost - - - - 10.06 2.48
One Model for All - - 22.50 19.20 9.12 1.49
TEMU-VTOFF 77.21 13.38 28.44 18.04 8.71 1.11
Results on the VITON-HD Dataset. In Table 2, we report the quantitative results on VITON-HD. In this setting, TEMU-VTOFF sets a new state-of-the-art across the majority of metrics, achieving the best scores for DISTS, FID, and KID. This indicates a superior ability to reconstruct structural details and to match the distribution of the ground-truth images. Notably, One Model for All achieves a competitive LPIPS score, which we cite from the original paper as public checkpoints are unavailable. Our method, however, achieves more robust performance on FID, KID, and DISTS. Since LPIPS and DISTS are critical in paired settings, we provide a qualitative analysis (Sec. E of the Appendix) showing LPIPS can diverge from human judgment, while DISTS aligns more reliably.
Overall, our method achieves solid improvements on VITON-HD, although the performance gains are less pronounced than on Dress Code. This is expected, as VITON-HD focuses exclusively on upper-body garments and is therefore a simpler benchmark. In contrast, the diverse and multi-category nature of Dress Code, with dresses, skirts, and pants, highlights the advantages of our approach, where the joint use of textual descriptions and fine-grained masks proves critical for accurate garment reconstruction. Accordingly, the strengths of our method are most evident in complex, multi-category scenarios. A visual comparison on sample VITON-HD images is shown in Fig. 4, which further demonstrates the improved garment reconstruction quality of our proposed method.
@@ 4.2 Ablation Studies
To assess the contribution of each component in our pipeline, we conduct a detailed ablation study on the Dress Code dataset reported in Table 3. We first investigate the impact of our dual-stream DiT architecture by removing the feature extractor $F_{E}$. In this setting, the garment aligner component is not employed. As shown, without the feature extractor, we experience a clear performance drop. In contrast, injecting $t=0$ keys and values from $F_{E}$ into the generator component through the proposed MHA operator enables richer, multi-scale conditioning, leading to better results. Then, we analyze the impact of employing the garment aligner module. As it can be seen, the aligner module helps to improve perceptual fidelity, particularly in categories with complex structures such as dresses, confirming that the designed components plays a critical role to the final performance.
Finally, removing garment descriptions or fine-grained masks consistently reduces performance, with the largest drop when both are absent, confirming that masks act as spatial anchors while text provides complementary semantic and category-level cues. The best results are obtained when both inputs are present, highlighting their complementarity.
@@
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ DISTS $\downarrow$ FID $\downarrow$
Effect of Dual-Stream DiT (w/o Garment Aligner)
w/o feature extractor $F_{E}$ 72.79 11.45 38.61 23.56 9.11 1.70 24.97 14.13 23.20 19.54 22.52 16.82
TEMU-VTOFF 76.01 12.85 30.84 20.63 5.91 0.78 21.77 11.26 22.26 14.22 17.86 11.86
Effect of Garment Aligner Component
w/o garment aligner 76.01 12.85 30.84 20.63 5.91 0.78 21.77 11.26 22.26 14.22 17.86 11.86
TEMU-VTOFF 75.95 12.90 31.46 18.66 5.74 0.65 19.75 10.94 19.57 13.83 16.67 11.29
@@
w/o text and masks 71.04 10.92 39.68 25.20 9.63 3.17 23.71 19.75 65.85 49.19 20.12 15.47
w/o text modulation 73.88 12.28 34.63 22.54 7.75 1.52 24.02 13.48 24.33 18.13 19.27 13.30
w/o fine-grained masks 74.65 12.30 32.33 20.87 6.58 1.03 20.85 11.31 22.34 15.74 19.42 13.62
TEMU-VTOFF 75.95 12.90 31.46 18.66 5.74 0.65 19.75 10.94 19.57 13.83 16.67 11.29
@@
Figure 5: Qualitative comparisons demonstrating the contribution of each proposed component. Left: effect of text and mask conditioning as “soft” and “hard” proxies for multi-category modeling. Right: impact of the lightweight garment aligner, which guides training with clean garment supervision.
To better understand the strength of each component proposed in our approach, we provide a visual comparison on Dress Code in Fig. 5. When our method relies exclusively on visual features from the person, without any textual guidance, it can struggle to resolve ambiguities in the garment design, leading to errors in structural elements such as neckline, sleeve length, or overall fit. The introduction of a textual description provides essential structural cues, enabling the model to capture the intended garment type and style. The fine-grained mask then imposes a precise spatial boundary, enforcing a clean silhouette and sharp edges, which improves the overall shape and contour of the garment. Finally, the garment aligner further improves the visual fidelity by encouraging the reconstruction of high-frequency details. This results in improved textures and more accurate patterns, ensuring that the final generated garment is not only structurally correct but also rich in fine-grained detail.
4.3 Cross-Dataset Generalization
To evaluate the robustness of TEMU-VTOFF against domain shifts and its ability to generalize to unseen garment types and poses, we conduct cross-dataset experiments. Specifically, we train our model on one dataset and evaluate it directly on the test set of the other dataset. We compare against MGT (Velioglu et al., 2025), TryOffDiff (Velioglu et al., 2024), and TryOffAnyone (Xarchakos and Koukopoulos, 2024). Note that we exclude Any2AnyTryon (Guo et al., 2025) from this specific analysis, as it is trained on a mixture of datasets including both Dress Code and VITON-HD, making the cross-dataset evaluation unfair. In Table 4, we present cross-dataset generalization results under two transfer settings. When trained on Dress Code and evaluated on VITON-HD, our model consistently surpasses MGT across both perceptual and distributional metrics, achieving a notably lower FID (20.39 vs. 23.11). Conversely, when trained on VITON-HD and tested on Dress Code (upper-body), our method again shows stronger generalization, obtaining an FID of 18.63 and clearly outperforming both TryOffDiff and TryOffAnyone.
Table 4: Quantitative comparison where models are trained on one dataset and tested on another to evaluate robustness to domain shift. $\uparrow$ indicates higher is better, $\downarrow$ lower is better.
Method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$
Training on Dress Code $\rightarrow$ Test on VITON-HD
MGT 74.26 10.24 42.57 28.73 23.11 10.81
TEMU-VTOFF 72.80 10.85 40.19 24.20 20.39 7.00
Training on VITON-HD $\rightarrow$ Test on Dress Code (Upper-Body)
TryOffDiff 75.33 11.50 44.64 32.14 41.91 21.78
TryOffAnyone 71.96 10.52 47.14 27.54 24.45 9.84
TEMU-VTOFF 73.36 11.51 39.74 23.84 18.63 6.31
Table 5: VTON results from CatVTON trained on two Dress Code variants: the original dataset and the version augmented with TEMU-VTOFF-generated images.
All Upper-Body
Training Dataset SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$
Dress Code 90.65 23.03 7.12 9.18 4.56 1.34 92.93 24.32 5.33 7.66 9.58 2.04
Dress Code (Augm.) 90.65 23.36 7.00 9.00 4.15 1.16 90.94 24.39 5.33 7.54 9.26 1.74
Lower-Body Dresses
Method SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ KID $\downarrow$
Dress Code 91.46 24.44 6.03 7.84 9.60 1.71 87.50 21.18 10.01 12.04 9.58 1.26
Dress Code (Augm.) 91.48 24.50 5.95 7.63 9.02 1.38 87.50 21.21 10.00 12.02 9.45 1.12
4.4 Downstream Utility
To demonstrate the practical utility of TEMU-VTOFF, we evaluate its effectiveness as a data augmentation tool for the VTON downstream task. High-quality paired data (i.e., person and in-shop garment) is expensive to acquire; our method addresses this by synthetically generating the “in-shop” garment directly from images of clothed people.
Experimental Setup. We use the Dress Code dataset (Morelli et al., 2022) and employ TEMU-VTOFF to generate synthetic in-shop garment images for the training samples. Specifically, for each person image in the upper- and lower-body categories, we generate the missing in-shop garment: the lower-body item for upper-body images and the upper-body item for lower-body images. This procedure augments the dataset with additional person-garment pairs generated by TEMU-VTOFF. We then employ CatVTON (Chong et al., 2025) (utilizing the SD 3 medium backbone) in two distinct settings: trained only on the standard Dress Code training set and trained on the Dress Code training set augmented with the synthetic pairs generated by TEMU-VTOFF.
Results. We evaluate the trained models on the official Dress Code test set. As shown in Table 5, the model trained with our augmented data achieves consistent improvements across both perceptual and distributional metrics. Notably, in the upper-body and lower-body categories, the augmented training yields lower FID scores (9.27 vs. 9.58 and 9.02 vs. 9.60, respectively). This confirms that the garments generated by TEMU-VTOFF preserve sufficient structural fidelity and texture details to serve as effective training signals, improving the generalization of state-of-the-art VTON models.